14 Check my work The following information applies to the questions displayed below Actuary and trustee reports indicate the following changes in the P8O and plan assets of Lakeside Cable during 2018: art 3 of 7 Prior service cost at Jan. 1, 2018, from plan anendment at the beginning of 2016 (amortization: $6 million per year) Net loss-pensions at Jan.1, 2018 (previous losses exceeded previous gains) Average remaining service life of the active employ Actuarys discount rate $40 million $58 nil1ion 10 years ee group ints 10 ($ in millions) eBook Print Plan $200 18 РВО Beginning of 2018 Service cost Interest cost, 108 Loss (gain) on PBO Less: Retiree benefits End of 2018 $380 Beginning of 20118 6 Return on plan assets 8 9.0 (108 expected) (2) cash contributions 42) Less: Retiree benefits 430 End of 2018 64 42) $240 Problem 17-16 Part 3 3. Prepare a pension spreadsheet to assist you in determining end of 2018 balances in the PBO, plan assets, prior service cost-AOCI, the net loss-AOCI, and the pension liability. (Enter your answers in millions (i.e.. 10,000,000 should be entered as 10). Enter credit amounts with a minus sign and debit amounts with a positive sign.) < Prey 14 15 16 18 of 20 Next >
Check my work Required information 3. Prepare a pension spreadsheet to assist you in determining end of 2018 balances in the PBO, plan assets, prior service cost-AOCI, the net loss-AOCI, and the pension liability, (Enter your answers in millions (i.e., 10,000,000 should be entered as 10). Enter credit amounts with a minus sign and debit amounts with a positive sign.) (S in millions) o Plan AssetsCost-AOCIA РВО Prior Service Net Loss Pension -AOCI Cash (Liability)/ Bal., Jan. 1, 2018 Service cost Interest cost, 10% Expected return on assets Loss on assets Amortization of Prior service cost-AOCI Net loss-AOCI Gain on PBO Cash contributions Retiree benefits Bal., Dec. 31, 2018 (380) (56) (38) 200 40 58 56 (56) (38) 64 (42 240 (64) 64 42 (430) 34 56 (10) Prex 14 151. 18f 20 Next>
Answer #1

PBO

Plan Assets

Prior service cost - AOCI

Net loss - AOCI

Pension expense

Cash

Net pension (liability)/Asset

Bal., Jan. 1, 2018

(380)

200

40

58

(180)

Service cost

(56)

56

(56)

Interest cost, 10%

(38)

38

(38)

Expected return on assets

20

(20)

20

Loss on assets

(2)

2

(2)

Amortization of:

Prior service cost – AOCI

(6)

6

Net loss – AOCI

(2)

2

Gain on PBO

2

(2)

2

Cash contributions

64

(64)

64

Retiree benefits

42

(42)

Bal., Dec. 31, 2018

(430)

240

34

56

82

(190)

380*10% = 38

200*10% = 20

Loss on assets = 20-18=2

Net loss – AOCI = (28-38)/10 years = 2